Suppose

A= | 4 3 5 |
| 1 3 - 5 |
| 2 1 5 |.

1) Evaluate det(A) by expanding along the second row. No Other Method.

2) Can cramers rule be used to solve the system :

A | x |. | 0 |
| y | = | 0 |
| z | | a |

Where A is given and a can't equal 0?

Give reasons. IF YES, solve with cramers rule. IF NOT, solve with other method.

Please assist.
1
Expert's answer
2020-04-27T16:05:44-0400

Given matrix A = "\left[ {\begin{array}{cc}
 4 & 3 & 5 \\
 1 & 3 & -5 \\
 2 & 1 & 5 \\
 \end{array} } \right]"

1) By expanding along the second row,

det(A) = (-1) "\begin{vmatrix}
 3 & 5 \\
 1 & 5
\end{vmatrix}" + 3 "\begin{vmatrix}
 4 & 5 \\
 2 & 5
\end{vmatrix}" - (-5) "\begin{vmatrix}
 4 & 3 \\
 2 & 1
\end{vmatrix}"

= (-1)(15-5) + 3(20-10) + 5(4-6) = -10 + 30 -10 = 10


2) Now since D = det(A) "\neq" 0, so Cramers rule can be used to solve the given system of equation AX=b where X = [x y z]' , b = [0 0 a]' and a "\neq" 0.

Now Dx = "\left|
\begin{array}{cc} 
0 & 3 & 5 \\ 
0 & 3 & -5 \\
a & 1 & 5 \\ 
\end{array}
\right|" = a(-15-15) = -30a "\neq" 0 (expanding along the first column)


Dy = "\left|
\begin{array}{cc} 
4 & 0 & 5 \\ 
1 & 0 & -5 \\
2 & a & 5 \\ 
\end{array}
\right|" = -a(-20-5) = 25a ​"\neq" 0 (expanding along the second column)


Dz = "\left|
\begin{array}{cc} 
4 & 3 & 0 \\ 
1 & 3 & 0 \\
2 & 1 & a \\ 
\end{array}
\right|" = a(12-3) = 9a "\neq" 0 (expanding along the third column)


So solution of the given system is

"x = \frac{D_x}{D} = \frac{-30a}{10} = -3a, y = \frac{D_y}{D} = \frac{25a}{10} = 2.5a, z = \frac{D_z}{D} = \frac{9a}{10} = 0.9a"